An electric bulb is marked 200 volt, 100 watt. Calculate resistance of its filament. If five such bulbs are connected in series to a 200 volt supply, how much current will flow through them?.

Grade:12

1 Answers

Satyam Kumar
6 years ago
W=V^2/R. R=V^2/W=(200*200)/100=400ohm. Req=5(400)=2000ohm. I=V/R=200/2000=1/10=0.1AMay this help you